# On the discrete spectrum of non-selfadjoint operators with applications to Schrödinger operators with complex potentials

### v1

@@ Abstract.
For relatively form-compact perturbations of non-negative selfadjoint operators, we obtain an upper bound on the number of discrete eigenvalues in half-planes separated from the positive real axis. The bound is given in terms of a partial trace of the real part of the Birman–Schwinger operator, or an appropriate rotation thereof. As an application to Schrödinger operators, we generalise the Cwikel–Lieb–Rozenblum inequality to complex potentials and derive new Lieb–Thirring type inequalities.
@@ 1. Introduction
$\gamma\geq 1/2\quad(d=1),\quad\gamma>0\quad(d=2),\quad\gamma\geq 0\quad(d=3),$ (2)
@@ 3. Applications to non-selfadjoint Schrödinger operators
$\gamma\geq 1/2\quad(d=1),\quad\gamma>0\quad(d=2),\quad\gamma\geq 0\quad(d=3).$
@@ 3.1. CLR inequalities for complex potentials
In the following result we derive bounds on $N(\mathop{\mathrm{Re}}\nolimits\lambda+\alpha\mathop{\mathrm{Im}}\nolimits\lambda<-\varepsilon;-\Delta+V)$ in terms of $\varepsilon$ and the $L^{\gamma+d/2}({\mathbb{R}}^{d})$-norm of $(\mathop{\mathrm{Re}}\nolimits V+\alpha\mathop{\mathrm{Im}}\nolimits V)_{-}$. In particular, the case $d\geq 3$, $\gamma=0$ and $\alpha=0$ is the generalisation of the CLR inequality to complex potentials.
@@
First we consider $p>1$ and set $q=2p$. To determine $\|u\|_{L^{q,\infty}({\mathbb{R}}^{d})}$ in (9), we shall note that $|u(\xi)|=\frac{1}{2\pi}(|\xi|^{2}+\varepsilon)^{-1/2}>t$ if and only if $|\xi|<((2\pi t)^{-2}-\varepsilon)^{1/2}$, which makes sense only for $t<\varepsilon^{-1/2}/(2\pi)$. Let $\tau_{d}$ denote the volume of the unit ball in ${\mathbb{R}}^{d}$. For $0<t<\varepsilon^{-1/2}/(2\pi)$ we define the function
@@ 3.2. LT inequalities for complex potentials
Let $p=d/2+\gamma$ with $\gamma>1/2$ for $d=1$ and $\gamma>0$ for $d\geq 2$. Then there exists a constant $C_{d,p}>0$ such that for all $V\in L^{p}({\mathbb{R}}^{d})$ and $\alpha\in{\mathbb{R}}$,
$\sum_{\lambda_{j}\in\sigma_{\rm d}(-\Delta+V)}(\mathop{\mathrm{Re}}\nolimits\lambda_{j}+\alpha\mathop{\mathrm{Im}}\nolimits\lambda_{j})_{-}^{\gamma}\leq C_{d,p}\,\int_{{\mathbb{R}}^{d}}(\mathop{\mathrm{Re}}\nolimits V(x)+\alpha\mathop{\mathrm{Im}}\nolimits V(x))_{-}^{p}\,{\rm d}x.$
@@
Let $\alpha\in{\mathbb{R}}$ and $\varepsilon>0$. Define $\varepsilon^{\prime}=2\varepsilon$ and $p^{\prime}=d/2+\gamma^{\prime}$ where $\gamma^{\prime}=1/2$ for $d=1$, $\gamma^{\prime}\in(0,\gamma)$ for $d=2$ and $\gamma^{\prime}=0$ for $d\geq 3$. Then $V\in L^{p^{\prime}}({\mathbb{R}}^{d})$. First note that
$N(\mathop{\mathrm{Re}}\nolimits\lambda+\alpha\mathop{\mathrm{Im}}\nolimits\lambda<-\varepsilon^{\prime};-\Delta+V)=N(\mathop{\mathrm{Re}}\nolimits\lambda+\alpha\mathop{\mathrm{Im}}\nolimits\lambda<-\varepsilon;-\Delta+V+\varepsilon).$
@@
$N(\mathop{\mathrm{Re}}\nolimits\lambda+\alpha\mathop{\mathrm{Im}}\nolimits\lambda<-\varepsilon;-\Delta+V+\varepsilon)$
@@
$N(\mathop{\mathrm{Re}}\nolimits\lambda+\alpha\mathop{\mathrm{Im}}\nolimits\lambda<-\varepsilon^{\prime};-\Delta+V)$ $=N(\mathop{\mathrm{Re}}\nolimits\lambda+\alpha\mathop{\mathrm{Im}}\nolimits\lambda<-\varepsilon;-\Delta+V+\varepsilon)$
$\leq C_{d,p^{\prime}}\,\varepsilon^{-\gamma^{\prime}}\int_{{\mathbb{R}}^{d}}(\mathop{\mathrm{Re}}\nolimits V(x)+\alpha\mathop{\mathrm{Im}}\nolimits V(x)+\varepsilon)_{-}^{p^{\prime}}\,{\rm d}x$
$=C_{d,p^{\prime}}\,\varepsilon^{-\gamma^{\prime}}\int_{{\mathbb{R}}^{d}}(\mathop{\mathrm{Re}}\nolimits V(x)+\alpha\mathop{\mathrm{Im}}\nolimits V(x)+\varepsilon)_{-}^{d/2+\gamma^{\prime}}\,{\rm d}x.$
Since $\gamma>\gamma^{\prime}$, an argument due to Aizenman and Lieb [1, Eq. (9)] yields, with substitution $\varepsilon^{\prime}=2\varepsilon$,
@@
$\leq C_{d,p^{\prime}}\,2^{\gamma}\,\gamma\int_{0}^{\infty}\varepsilon^{\gamma-\gamma^{\prime}-1}\int_{{\mathbb{R}}^{d}}(\mathop{\mathrm{Re}}\nolimits V(x)+\alpha\mathop{\mathrm{Im}}\nolimits V(x)+\varepsilon)_{-}^{d/2+\gamma^{\prime}}\,{\rm d}x\,{\rm d}\varepsilon$
$=C_{d,p^{\prime}}\,2^{\gamma}\gamma\,\frac{\Gamma(d/2+\gamma^{\prime}+1)\Gamma(\gamma-\gamma^{\prime})}{\Gamma(d/2+\gamma+1)}\int_{{\mathbb{R}}^{d}}(\mathop{\mathrm{Re}}\nolimits V(x)+\alpha\mathop{\mathrm{Im}}\nolimits V(x))_{-}^{d/2+\gamma}\,{\rm d}x.$
@@ Acknowledgements
The PhD of S. P. is funded through a Development and Promotion of Science and Technology (DPST) scholarship of the Royal Thai Government, ref. 5304.1/3758. The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Geometric spectral theory and applications, where work on this paper was undertaken. This work was supported by EPSRC grant EP/Z000580/1.

### v2

@@ Abstract.
For relatively form-compact perturbations of non-negative selfadjoint operators, we obtain an upper bound on the number of discrete eigenvalues in half-planes separated from the positive real axis. The bound is given in terms of a partial trace of the real part of the Birman–Schwinger operator, or an appropriate rotation thereof. While eigenvalue counting estimates of this type are classical in the selfadjoint setting, no analogous connection between the number of discrete eigenvalues and the Birman–Schwinger operator has previously been established in the non-selfadjoint theory. The proof proceeds via techniques in antisymmetric tensor product spaces that serve as a non-selfadjoint replacement for the classical arguments. As an application to Schrödinger operators, we generalise the Cwikel–Lieb–Rozenblum inequality to complex potentials and derive new Lieb–Thirring type inequalities. We also analyse the sharpness of the obtained bounds and discuss their optimality within the considered framework.
@@ 1. Introduction
$\gamma\geq 1/2\quad(d=1),\quad\gamma>0\quad(d=2),\quad\gamma\geq 0\quad(d\geq 3),$ (2)
@@ 3. Applications to non-selfadjoint Schrödinger operators
$\gamma\geq 1/2\quad(d=1),\quad\gamma>0\quad(d=2),\quad\gamma\geq 0\quad(d\geq 3).$
@@ 3.1. CLR inequalities for complex potentials
In the following result we derive bounds on $N(\mathop{\mathrm{Re}}\nolimits\lambda+\alpha\mathop{\mathrm{Im}}\nolimits\lambda<-\varepsilon;-\Delta+V)$ in terms of $\varepsilon$ and the $L^{\gamma+d/2}({\mathbb{R}}^{d})$-norm of $(\mathop{\mathrm{Re}}\nolimits V+\alpha\mathop{\mathrm{Im}}\nolimits V)_{-}$. In particular, the case $d\geq 3$ and $\gamma=0$ is the generalisation of the CLR inequality to complex potentials. The inequality for general $d$ and $\gamma$ is the generalisation of e.g. [28, Eq. (2.8)].
@@
The constants can be taken as
$C_{d,p}=\begin{cases}1/2&\text{if }d=1,\,p=1,\\ 2^{p+\gamma-4}\pi^{-2p}\tau_{d}\,d^{d/2}\gamma^{\gamma}p^{p+1}(p-1)^{1-2p}&\text{if }p>\max\{1,d/2\},\\ 2^{-5}\pi^{-d}\tau_{d}\,d^{d+1}(d/2-1)^{1-d}&\text{if }d\geq 3,\,p=d/2,\end{cases}$
where $\tau_{d}$ denotes the volume of the unit ball in ${\mathbb{R}}^{d}$.
@@
First we consider $p>1$ and set $q=2p$. To determine $\|u\|_{L^{q,\infty}({\mathbb{R}}^{d})}$ in (9), we shall note that $|u(\xi)|=\frac{1}{2\pi}(|\xi|^{2}+\varepsilon)^{-1/2}>t$ if and only if $|\xi|<((2\pi t)^{-2}-\varepsilon)^{1/2}$, which makes sense only for $t<\varepsilon^{-1/2}/(2\pi)$. Recall that $\tau_{d}$ denotes the volume of the unit ball in ${\mathbb{R}}^{d}$. For $0<t<\varepsilon^{-1/2}/(2\pi)$ we define the function
@@
Remark 10.
Using the estimates in Theorem 8 on the number of discrete eigenvalues for a given parameter $\gamma$, in Theorem 13 below we prove LT inequalities for all parameters that are larger than $\gamma$. On the other hand, once we know that an LT inequality of the form
$\sum_{\lambda_{j}\in\sigma_{\rm d}(-\Delta+V)}(\mathop{\mathrm{Re}}\nolimits\lambda_{j}+\alpha\mathop{\mathrm{Im}}\nolimits\lambda_{j})_{-}^{\gamma}\leq\widetilde{C}_{d,p}\,\int_{{\mathbb{R}}^{d}}(\mathop{\mathrm{Re}}\nolimits V(x)+\alpha\mathop{\mathrm{Im}}\nolimits V(x))_{-}^{p}\,{\rm d}x$
holds with a constant $\widetilde{C}_{d,p}>0$ that is independent of $\alpha$ and $V$, we obtain an estimate on the number of discrete eigenvalues as follows: using the lower bound
$\sum_{\lambda_{j}\in\sigma_{\rm d}(-\Delta+V)}(\mathop{\mathrm{Re}}\nolimits\lambda_{j}+\alpha\mathop{\mathrm{Im}}\nolimits\lambda_{j})_{-}^{\gamma}\geq\sum_{\lambda_{j}\in\sigma_{\rm d}(-\Delta+V),\atop\mathop{\mathrm{Re}}\nolimits\lambda_{j}+\alpha\mathop{\mathrm{Im}}\nolimits\lambda_{j}<-\varepsilon}\varepsilon^{\gamma}=\varepsilon^{\gamma}N(\mathop{\mathrm{Re}}\nolimits\lambda+\alpha\mathop{\mathrm{Im}}\nolimits\lambda<-\varepsilon;-\Delta+V),$
we obtain
$N(\mathop{\mathrm{Re}}\nolimits\lambda+\alpha\mathop{\mathrm{Im}}\nolimits\lambda<-\varepsilon;-\Delta+V)\leq\widetilde{C}_{d,p}\,\varepsilon^{-\gamma}\int_{{\mathbb{R}}^{d}}(\mathop{\mathrm{Re}}\nolimits V(x)+\alpha\mathop{\mathrm{Im}}\nolimits V(x))_{-}^{p}\,{\rm d}x.$
In particular, if $C_{d,p}^{\rm\,sharp}$ denotes the sharp constant in Theorem 8 and $\widetilde{C}_{d,p}^{\rm\,sharp}$ denotes the sharp constant in Theorem 13, the above argument implies that
$C_{d,p}^{\rm\,sharp}\leq\widetilde{C}_{d,p}^{\rm\,sharp}.$
In the endpoint case in one dimension, our constant is sharp.
Theorem 11.
Let $d=1$ and $p=1$. Then the constant $C_{d,p}=1/2$ in Theorem 8 is sharp.
Proof.
Let $V_{t}=-t\delta$ for $t>0$ where $\delta\in L^{1}({\mathbb{R}})$ is the delta-potential. Since $V_{t}$ is real-valued, the spectrum is real and the inequality in Theorem 8 is independent of $\alpha$. Thus it suffices to prove optimality for $\alpha=0$. For any $t>0$ the point $\lambda_{t}=-t^{2}/4$ is a discrete eigenvalue with corresponding eigenfunction $\psi_{t}(x)=\exp(-t|x|/2)$; in fact, this is the only discrete eigenvalue. In order that $\lambda_{t}<-\varepsilon$, we require $t>2\varepsilon^{1/2}$. Then
$\frac{N(\lambda<-\varepsilon;-\frac{\rm d^{2}}{{\rm d}x^{2}}+V_{t})}{\varepsilon^{-\gamma}\,\int_{{\mathbb{R}}}(V_{t}(x))_{-}\,{\rm d}x}=\frac{1}{\varepsilon^{-\gamma}t}=\frac{1}{\varepsilon^{-1/2}t}$
where we have used that $\gamma=1/2$. Now the claim follows since, in the limit as $t\to 2\varepsilon^{1/2}$, the right-hand side converges to $1/2=C_{d,p}$. ∎
Remark 12.
For other combinations of $d$ and $p$ the sharpness of the constant $C_{d,p}$ in Theorem 8 is an open problem. For $d\geq 3$ and $\gamma=0$, the constant $C_{d,d/2}$ is the same as Cwikel’s constant in his proof of the CLR inequality, see [11]. In the selfadjoint case, the sharp constant is unknown as well but it is known that e.g. Lieb’s proof [30] gives a constant better than Cwikel’s constant. The sharpness of the constants in the non-selfadjoint LT inequalities (see Theorem 13 below) is also an open problem. For an overview of the results working towards sharp constants in selfadjoint CLR and LT inequalities, see [17]. It is important to remark that, since here we allow for complex potentials as well, the sharp constants may be larger compared to the selfadjoint case.
@@ 3.2. LT inequalities for complex potentials
Let $p=d/2+\gamma$ with $\gamma>1/2$ for $d=1$ and $\gamma>0$ for $d\geq 2$. Then there exists a constant $\widetilde{C}_{d,p}>0$ such that for all $V\in L^{p}({\mathbb{R}}^{d})$ and $\alpha\in{\mathbb{R}}$,
$\sum_{\lambda_{j}\in\sigma_{\rm d}(-\Delta+V)}(\mathop{\mathrm{Re}}\nolimits\lambda_{j}+\alpha\mathop{\mathrm{Im}}\nolimits\lambda_{j})_{-}^{\gamma}\leq\widetilde{C}_{d,p}\,\int_{{\mathbb{R}}^{d}}(\mathop{\mathrm{Re}}\nolimits V(x)+\alpha\mathop{\mathrm{Im}}\nolimits V(x))_{-}^{p}\,{\rm d}x.$
The constant can be taken as
$\widetilde{C}_{d,p}=C_{d,p^{\prime}}^{\,\rm sharp}\,\frac{\gamma^{\gamma+1}}{(\gamma^{\prime})^{\gamma^{\prime}}(\gamma-\gamma^{\prime})^{\gamma-\gamma^{\prime}}}\,\frac{\Gamma(d/2+\gamma^{\prime}+1)\Gamma(\gamma-\gamma^{\prime})}{\Gamma(d/2+\gamma+1)}$ (12)
where $C_{d,p^{\prime}}^{\,\rm sharp}$ is the sharp constant in Theorem 8 for $p^{\prime}=d/2+\gamma^{\prime}$ with
$\gamma^{\prime}\in[1/2,\gamma)\quad(d=1),\quad\gamma^{\prime}\in(0,\gamma)\quad(d=2),\quad\gamma^{\prime}\in[0,\gamma)\quad(d\geq 3);$ (13)
we employ the convention that $(\gamma^{\prime})^{\gamma^{\prime}}=1$ if $\gamma^{\prime}=0$.
@@
Let $\alpha\in{\mathbb{R}}$ and $\varepsilon>0$. Define $\varepsilon^{\prime}=(1+t)\varepsilon$ for some $t>0$, and let $p^{\prime}=d/2+\gamma^{\prime}$ where $\gamma^{\prime}$ is as in (13). Then $V\in L^{p^{\prime}}({\mathbb{R}}^{d})$. First note that
$N(\mathop{\mathrm{Re}}\nolimits\lambda+\alpha\mathop{\mathrm{Im}}\nolimits\lambda<-\varepsilon^{\prime};-\Delta+V)=N(\mathop{\mathrm{Re}}\nolimits\lambda+\alpha\mathop{\mathrm{Im}}\nolimits\lambda<-t\varepsilon;-\Delta+V+\varepsilon).$
@@
$N(\mathop{\mathrm{Re}}\nolimits\lambda+\alpha\mathop{\mathrm{Im}}\nolimits\lambda<-t\varepsilon;-\Delta+V+\varepsilon)$
@@
$N(\mathop{\mathrm{Re}}\nolimits\lambda+\alpha\mathop{\mathrm{Im}}\nolimits\lambda<-\varepsilon^{\prime};-\Delta+V)$ $=N(\mathop{\mathrm{Re}}\nolimits\lambda+\alpha\mathop{\mathrm{Im}}\nolimits\lambda<-t\varepsilon;-\Delta+V+\varepsilon)$
$\leq C_{d,p^{\prime}}^{\,\rm sharp}\,(t\varepsilon)^{-\gamma^{\prime}}\int_{{\mathbb{R}}^{d}}(\mathop{\mathrm{Re}}\nolimits V(x)+\alpha\mathop{\mathrm{Im}}\nolimits V(x)+\varepsilon)_{-}^{p^{\prime}}\,{\rm d}x$
$=C_{d,p^{\prime}}^{\,\rm sharp}\,t^{-\gamma^{\prime}}\varepsilon^{-\gamma^{\prime}}\int_{{\mathbb{R}}^{d}}(\mathop{\mathrm{Re}}\nolimits V(x)+\alpha\mathop{\mathrm{Im}}\nolimits V(x)+\varepsilon)_{-}^{d/2+\gamma^{\prime}}\,{\rm d}x.$
Since $\gamma>\gamma^{\prime}$, an argument due to Aizenman and Lieb [1, Eq. (9)] yields, with substitution $\varepsilon^{\prime}=(1+t)\varepsilon$,
@@
$\leq C_{d,p^{\prime}}^{\,\rm sharp}\,t^{-\gamma^{\prime}}(1+t)^{\gamma}\,\gamma\int_{0}^{\infty}\varepsilon^{\gamma-\gamma^{\prime}-1}\int_{{\mathbb{R}}^{d}}(\mathop{\mathrm{Re}}\nolimits V(x)+\alpha\mathop{\mathrm{Im}}\nolimits V(x)+\varepsilon)_{-}^{d/2+\gamma^{\prime}}\,{\rm d}x\,{\rm d}\varepsilon$
$=C_{d,p^{\prime}}^{\,\rm sharp}\,t^{-\gamma^{\prime}}(1+t)^{\gamma}\,\gamma\,\frac{\Gamma(d/2+\gamma^{\prime}+1)\Gamma(\gamma-\gamma^{\prime})}{\Gamma(d/2+\gamma+1)}\int_{{\mathbb{R}}^{d}}(\mathop{\mathrm{Re}}\nolimits V(x)+\alpha\mathop{\mathrm{Im}}\nolimits V(x))_{-}^{d/2+\gamma}\,{\rm d}x.$
An easy computation shows that the infimum over $t>0$ is attained at $t=\frac{\gamma^{\prime}}{\gamma-\gamma^{\prime}}$ with
$\inf_{t>0}\,t^{-\gamma^{\prime}}(1+t)^{\gamma}=\frac{\gamma^{\gamma}}{(\gamma^{\prime})^{\gamma^{\prime}}(\gamma-\gamma^{\prime})^{\gamma-\gamma^{\prime}}}.$
@@
Remark 15.
For $\gamma\geq 1$ it was shown in [16] that an inequality as in Theorem 13 holds and the constant $\widetilde{C}_{d,p}$ can be taken as the sharp selfadjoint LT constant, which is therefore also the sharp constant here. For $\gamma<1$ the sharp constant is unknown; whether it coincides with our $\widetilde{C}_{d,p}$ in (12) is left as an open problem. For $d\geq 3$, if we calculate $\widetilde{C}_{d,p}$ in (12) with $\gamma^{\prime}=0$ and use that $\gamma\,\Gamma(\gamma)=\Gamma(\gamma+1)$, we arrive at
$\widetilde{C}_{d,p}=C_{d,d/2}^{\,\rm sharp}\,\frac{\Gamma(d/2+1)\,\Gamma(\gamma+1)}{\Gamma(d/2+\gamma+1)}.$
Note that, with $\Gamma(1)=1$, we obtain $\widetilde{C}_{d,p}\to C_{d,d/2}^{\,\rm sharp}$ as $p\to d/2$, i.e. we recover the sharp CLR constant $C_{d,d/2}^{\,\rm sharp}$ from Theorem 8. Furthermore, we make the observation that $\widetilde{C}_{d,p}=C_{d}\,L_{\gamma,d}^{\mathrm{cl}}$ where
$C_{d}:=C_{d,d/2}^{\,\rm sharp}\,\Gamma(d/2+1)(4\pi)^{d/2},\quad L_{\gamma,d}^{\mathrm{cl}}:=(4\pi)^{-d/2}\,\frac{\Gamma(\gamma+1)}{\Gamma(d/2+\gamma+1)}.$
The semiclassical LT constant $L_{\gamma,d}^{\mathrm{cl}}$ is known to be a lower bound for the sharp selfadjoint LT constant, and hence for the sharp constant $\widetilde{C}_{d,p}^{\,\rm sharp}$ in Theorem 13. This implies the two-sided bound
$L_{\gamma,d}^{\mathrm{cl}}\leq\widetilde{C}_{d,p}^{\,\rm sharp}\leq C_{d}\,L_{\gamma,d}^{\mathrm{cl}}.$
@@ Acknowledgements
The authors are grateful to Lukas Schimmer for drawing their attention to reference [28]. The PhD of S. P. is funded through a Development and Promotion of Science and Technology (DPST) scholarship of the Royal Thai Government, ref. 5304.1/3758. The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Geometric spectral theory and applications, where work on this paper was undertaken. This work was supported by EPSRC grant EP/Z000580/1.